OpenStudy (mackenzie2013):

What is the value of the 24th term in the following arithmetic sequence? -9, -2, 5, 12, ... 152 -170 170 -200 Please Explain how to get the answer! :)

OpenStudy (lukebluefive):

If you look at each of the four given numbers in the sequence as the following pairs, a pattern emerges: -9, -2 -2, 5 5, 12 Can you see something similar among these three pairs?

OpenStudy (mackenzie2013):

Yes, I do.

OpenStudy (lukebluefive):

Well, if you've identified the pattern, then continue it, assuming that -9 is the first term in the sequence, and write out the sequence to the 24th term. That's your answer.

OpenStudy (lukebluefive):

So, whatever you do to -9 to get -2, -2 to get 5, and 5 to get 12, repeat on 12 to get the next term, then on that term to get the term after that, and so on.

OpenStudy (mackenzie2013):

But im given a formula to find out the answer.

OpenStudy (mackenzie2013):

an = a1 + d(n - 1)

OpenStudy (lukebluefive):

Okay, that's more complicated than it needs to be, but it works. an = The value of the nth term a1 = The first term in the sequence d = The difference between each successive term in the sequence n = The term number

OpenStudy (mackenzie2013):

okay I understand that.

OpenStudy (lukebluefive):

So now you can plug in the values for a1, d, and n to get the answer. Since you want the value of the 24th term, n=24.

OpenStudy (mackenzie2013):

a1= 9 + 7(n-1) but n is 24 so its A1=9+161 Is that right?

OpenStudy (lukebluefive):

a1 is actually -9

OpenStudy (mackenzie2013):

Whoops, my mistake.

OpenStudy (lukebluefive):

You're very close. The variable on the left side of the formula is actually an.

OpenStudy (mackenzie2013):

Whoops, Haha I'm new to this. If you can't tell.

OpenStudy (lukebluefive):

You're actually doing better than most, but everyone learns at a different rate.

OpenStudy (mackenzie2013):

The answer is 152?

OpenStudy (lukebluefive):

Yep, nicely done.

OpenStudy (mackenzie2013):

Thank you! :)

OpenStudy (lukebluefive):

You're very welcome! It was a pleasure helping you. Let me know if you have any other questions.